In the circuit below, the voltage supplied by the power source is 10 V and the current flowing through the circuit is 0.12 A.KE3 Prelaboratory Questions E3 - Question 1: Resistance in a Circuit 1 of 5 Review 1 Constants i Periodic Table Part A In the circuit below, the voltage supplied by the power source is 10 V and the current flowing through the circuit is 0.12 A What is the resistance R2? 35 S2 33.75 Submit Previous Answers Request Answer 10V X Incorrect; Try Again; 3 attempts remaining Provide Feedback Next >
Answer #1

Req = 35+R2

Also Req = V/I = 10/0.12 = 83.34 ohm

Hence R2 = 83.34 - 35 = 48.34 ohm